hsdfhsfd:

Sherry uses the steps below to solve the equation x + (negative 8) = 3 x + 6. Step 1 Add 1 negative x-tile to both sides and create zero pairs Step 2 Add 8 positive unit tiles to both sides and create zero pairs. Step 3 Divide the 14 unit tiles evenly among the 2 x-tiles. Step 4 The solution is x = 7 What is Sherry’s error? In step 1, she should have added 3 negative x-tiles to both sides. In step 2, she should have added 6 negative unit tiles to both sides. In step 3, she did not divide the tiles evenly. In step 4, she should have found the solution x = negative 1.

hsdfhsfd:

pleasw help me

hsdfhsfd:

ok im sorry my grammer isint the best

Hoodmemes:

Ight ima do this the right way

Hoodmemes:

Step one: \[x + (-8) + (-x)= 3x + 6 + (-x)\] \[- 8 = 2x +6\]

Hoodmemes:

Step Two: \[-8 + 8= 2x + 6 + 8\] \[0= 2x+ 14 \]

Hoodmemes:

Step Three: \[0 + (-14) = 2x + 14 + (-14)\] \[-14=2x\]

Hoodmemes:

Step Four: \[\frac{ -14 }{ 2 }= -7\]

Hoodmemes:

Step Five: \[x=(-blank)\]

Hoodmemes:

hope this helps